# Spatio-temporal methods for estimating subsurface ocean thermal response to tropical cyclones

### v3

@@ 5.6 Vertically averaged temperature differences
6 Discussion
In this paper, we have introduced a comprehensive methodological framework based upon functional ANOVA for analyzing tropical cyclone-induced ocean temperature changes. This framework includes a pairing process for identifying reference observations, the construction of an annual mean field to account for seasonal shift at the time scales of the measured differences, as well as estimation of observation variability and computationally efficient smoothers that leverage these estimates of variability.
@@ 6 Discussion
Of course, this is not the full scientific story. The ocean response to the passage of TCs is complex, including contributions from upwelling near the storm center, downwelling in the outer regions, and mixed layer entrainment (Elsberry, Fraim and Trapnell, 1976; Lin et al., 2017); horizontal currents (D’Asaro et al., 2007); in addition to air-sea exchanges of heat (Emanuel, 1986, 1999). The initial stratification in a region is also important for understanding the relevant processes. Nonetheless, the statistical methodology presented in this paper provides an important new framework through which TC-induced temperature changes in the upper ocean may be studied as a function of depth, distance from TC track center, and time since TC passage.
Although our methodological framework was developed with ocean thermal response in mind, the techniques discussed are of general interest. An immediate generalization is the use of the exact same methodology for Argo salinity profiles in lieu of temperature profiles. One may expect that due to rainfall, evaporation, and mixing, the passage of a tropical cyclone would be registered in the salinity profiles of the affected ocean region. Similar analyses for other transient atmospheric phenomena, including atmospheric rivers and southern ocean storm systems, are also possible. One could additionally regress these signals on latitude, storm strength, pre-storm ocean state, ocean basin, and other covariates to better model and understand the underlying scientific phenomena. Regression on latitude may be particularly relevant with regard to the ongoing scientific discourse on the large-scale climatological effects of tropical cyclones (Emanuel, 2001; Korty, Emanuel and Scott, 2008; Jansen and Ferrari, 2009; Jansen, Ferrari and Mooring, 2010; Haney et al., 2012).
@@
An additional extension of statistical and scientific interest involves carefully accounting for float drift. In this paper, we require the baseline and signal profiles to be within $0.2^{\circ}$ of each other and then assume that they share a single location in the analysis, admitting a useful simplification of the fitted covariance function. An extension would involve relaxing the strict closeness requirement of the paired profiles and removing the assumption that the paired profiles are in the same location by more carefully modeling their covariance, taking into account their relative spatial locations. This requires a nontrivial generalization of the Gaussian process model of Kuusela and Stein (2018), but would provide the benefit of increasing the number of available pairs with which to fit the model. Yet another extension would be the consideration of parametric models for the TC wake (Haney et al., 2012). Finally, the fusion of microwave sea surface temperature (mSST; Gentemann et al., 2003; Wentz, Gentemann and Hilburn, 2005) observations with Argo profiles might provide an opportunity to produce estimates on the individual-storm level, rather than in terms of the global aggregates, as was done here. mSST observations are very dense in latitude and longitude, but lack the depth penetration provided by the Argo profiles, which are relatively sparse in latitude and longitude. Therefore, mSST and Argo are naturally complementary in the information provided.
@@ Appendix A Derivation of the thin plate spline estimator
In this section, we motivate and derive the thin plate spline estimator (Duchon, 1977; Wahba, 1980, 1990; Green and Silverman, 1994; Wood, 2003) used in Section 4.5 for completeness. First, we review the natural thin plate spline smoother in Section A.1. Sections A.2 and A.3 then describe versions of the thin plate spline when subject to covariance reweighing (via generalized least squares; Rao, 1973; Draper and Smith, 1998; Nychka, 2000; Ruppert, Wand and Carroll, 2003) and when its basis functions are constructed from a fixed knot set (Wahba, 1980; Wood, 2017).
@@ A.2 The covariance-reweighed thin plate spline smoother
Due to the correlated and heteroskedastic nature of the temperature differences in this paper, we desire a version of the thin plate spline smoother (37) that properly accounts for the these effects. This is done via generalized least squares (Rao, 1973; Draper and Smith, 1998; Nychka, 2000; Ruppert, Wand and Carroll, 2003), which poses the objective function:
@@ A.3 Adaptation to fixed knots
Finally, we detail the adaptation of the covariance-reweighed thin plate spline smoother to the fixed-knot regime (Wahba, 1980, 1990; Ruppert, Wand and Carroll, 2003; Wood, 2017). Suppose a set of fixed knots $\tilde{\boldsymbol{\xi}}_{1},\dotsc,\tilde{\boldsymbol{\xi}}_{m}$, with $\tilde{\mathbf{R}}$ as defined in (23) and $\mathbf{S}$} as defined in (25). The thin plate spline smoothing problem, restricted to this set of knots, is given (Wood, 2017) by:
@@ Appendix B Leave-one-out cross-validation
(a) Leave-one-out cross-validation errors
(b) Chosen $\lambda$ values
@@
Acknowledgments
The first author acknowledges support from the NSF GRFP (Award DGE175016) and NSF DMS (Award 1520786). The fourth author was supported by NSF EarthCube (Award 1928305). The authors thank members of the CMU Statistical Methods for the Physical Sciences (STAMPS) Research Group and the SAMSI/CMU Statistical Oceanography Working Group for insightful comments and feedback during this work.
The Argo data used in this paper were collected and are made freely available by the International Argo Program and the national programs that contribute to it (http://doi.org/10.17882/42182). The Argo Program is part of the Global Ocean Observing System.
{supplement}
Supplement to “Spatio-temporal methods for estimating subsurface ocean thermal response to tropical cyclones” (DOI: XXXX; .pdf). We provide a comprehensive set of figures illustrating the coefficients and fitted values for the seasonal mean field, Gaussian process, and thin plate spline models in the supplementary document. The supplement, along with code implementing our methodology, may be found at https://github.com/huisaddison/tc-ocean-methods.

### v4

@@ 4.6 Vertically averaged temperature differences
[…]
[Discussion] In this paper, we have introduced a comprehensive methodological framework based upon an ANOVA-type decomposition for analyzing tropical cyclone-induced ocean temperature changes. This framework includes a pairing process for identifying reference observations, the construction of an annual mean field to account for seasonal shift at the time scales of the measured differences, as well as estimation of observation variability and computationally efficient smoothers that leverage these estimates of variability.
@@
Of course, this is not the full scientific story. The ocean response to the passage of TCs is complex, including contributions from upwelling near the storm center, downwelling in the outer regions, and mixed layer entrainment (Elsberry et al., 1976; Lin et al., 2017); horizontal currents (D’Asaro et al., 2007); in addition to air-sea exchanges of heat (Emanuel, 1986, 1999). The initial stratification in a region is also important for understanding the relevant processes. Nonetheless, the statistical methodology presented in this paper provides an important new framework through which TC-induced temperature changes in the upper ocean may be studied as a function of depth, distance from TC track center, and time since TC passage.
Although our methodological framework was developed with ocean thermal response in mind, the techniques discussed are of general interest. An immediate generalization is the use of the exact same methodology for Argo salinity profiles in lieu of temperature profiles. One may expect that due to rainfall, evaporation, and mixing, the passage of a tropical cyclone would be registered in the salinity profiles of the affected ocean region. Similar analyses for other transient atmospheric phenomena, including atmospheric rivers and southern ocean storm systems, are also possible. One could additionally regress these signals on latitude, storm strength, pre-storm ocean state, ocean basin, and other covariates to better model and understand the underlying scientific phenomena. Regression on latitude may be particularly relevant with regard to the ongoing scientific discourse on the large-scale climatological effects of tropical cyclones (Emanuel, 2001; Korty et al., 2008; Jansen and Ferrari, 2009; Jansen et al., 2010; Haney et al., 2012).
@@
An additional extension of statistical and scientific interest involves carefully accounting for float drift. In this paper, we require the baseline and signal profiles to be within $0.2^{\circ}$ of each other and then assume that they share a single location in the analysis, admitting a useful simplification of the fitted covariance function. An extension would involve relaxing the strict closeness requirement of the paired profiles and removing the assumption that the paired profiles are in the same location by more carefully modeling their covariance, taking into account their relative spatial locations. This requires a nontrivial generalization of the Gaussian process model of Kuusela and Stein (2018), but would provide the benefit of increasing the number of available pairs with which to fit the model. Yet another extension would be the consideration of parametric models for the TC wake (Haney et al., 2012). Finally, the fusion of microwave sea surface temperature (mSST; Gentemann et al., 2003; Wentz et al., 2005) observations with Argo profiles might provide an opportunity to produce estimates on the individual-storm level, rather than in terms of the global aggregates, as was done here. mSST observations are very dense in latitude and longitude, but lack the depth penetration provided by the Argo profiles, which are relatively sparse in latitude and longitude. Therefore, mSST and Argo are naturally complementary in the information provided.
[…]
Code implementing our methodology may be found at https://github.com/huisaddison/tc-ocean-methods.
@@ Appendix A Derivation of the thin plate spline estimator
In this section, provided for completeness, we motivate and derive the thin plate spline estimator (Duchon, 1977; Wahba, 1980, 1990; Green and Silverman, 1994; Wood, 2003) used in Section 3.5 for completeness. First, we review the natural thin plate spline smoother in Section A.1. Sections A.2 and A.3 then describe versions of the thin plate spline when subject to covariance reweighing (via generalized least squares; Rao, 1973; Draper and Smith, 1998; Nychka, 2000; Ruppert et al., 2003) and when its basis functions are constructed from a fixed knot set (Wahba, 1980; Wood, 2017).
@@ A.2 The covariance-reweighed thin plate spline smoother
Due to the correlated and heteroskedastic nature of the temperature differences in this paper, we desire a version of the thin plate spline smoother (37) that properly accounts for the these effects. This is done via generalized least squares (Rao, 1973; Draper and Smith, 1998; Nychka, 2000; Ruppert et al., 2003), which poses the objective function:
@@ A.3 Adaptation to fixed knots
Finally, we detail the adaptation of the covariance-reweighed thin plate spline smoother to the fixed-knot regime (Wahba, 1980, 1990; Ruppert et al., 2003; Wood, 2017). Suppose a set of fixed knots $\tilde{\boldsymbol{\xi}}_{1},\dotsc,\tilde{\boldsymbol{\xi}}_{m}$, with $\tilde{\mathbf{R}}$ as defined in (23) and $\mathbf{S}$} as defined in (25). The thin plate spline smoothing problem, restricted to this set of knots, is given (Wood, 2017) by:
@@ Appendix B Leave-one-out cross-validation
((a)) Leave-one-out cross-validation errors
((b)) Chosen $\lambda$ values
@@
[…]
AJH, MK, and ABL developed the methodology with domain knowledge and conceptualization from DG and KMW. AJH performed the formal analysis, visualization, and writing, with feedback from all authors. Software was developed by AJH extending original software by MK.
[…]
The authors declare that they have no conflict of interest.
Acknowledgements.
AJH acknowledges support from the NSF GRFP (Award DGE175016) and NSF DMS (Award 1520786). MK was supported by NOAA (Award NA21OAR4310258). DG was supported by NSF EarthCube (Award 1928305) and NOAA (Award NA21OAR4310261). The authors thank members of the CMU Statistical Methods for the Physical Sciences (STAMPS) Research Group and the SAMSI/CMU Statistical Oceanography Working Group for insightful comments and feedback during this work. The Argo data used in this paper were collected and are made freely available by the International Argo Program and the national programs that contribute to it (http://doi.org/10.17882/42182). The Argo Program is part of the Global Ocean Observing System.